OpenStudy (anonymous):

Find the diagonal of a cube if its side equals 5.

OpenStudy (anonymous):

\( 5 \times \sqrt{3} \)

OpenStudy (anonymous):

The longest diagonal of the cube!

OpenStudy (anonymous):

5^2+5^2=2*25=50=5sqrt(2) 25+5sqrt(2)=3*5^2=75=5sqrt(3)

OpenStudy (vishal_kothari):

\[3\sqrt{5}\]

OpenStudy (anonymous):

The answer is 5*sqrt(3), really we are, if we draw the cube, applying pythagorean theorm twice

OpenStudy (vishal_kothari):

8.66

OpenStudy (anonymous):

you are ?

OpenStudy (anonymous):

@Lagrange

OpenStudy (anonymous):

yeah@FoolForMath

OpenStudy (anonymous):

you dont believe me

OpenStudy (anonymous):

I am incredulous without proof :P

OpenStudy (anonymous):



OpenStudy (anonymous):

lol, I know the proof.

OpenStudy (anonymous):

Haven't I posted the same answer before ? :)

OpenStudy (anonymous):

The line from D to B in the diagram completes the triangle DCB, which has a right angle at C. Thus, by Pythagoras' Theorem: (DB)^2=(DC)^2+(CB)^2=5^2+5^2=50=5sqrt(2)

OpenStudy (anonymous):

Now consider the triangle ADB